Suppose the demand equation for computers by teetan ltd for the year 2017 is given by Qd=1200-P and the supply equation is given by Q=120+3P. Find equilibrium price and analyse what would be the excess demand or supply if price changes to rs 400 and rs 120
1
Expert's answer
2020-11-24T06:24:28-0500

At equilibrium price, the quantity demanded is equals to the quantity supplied to the market.

This implies that Demand=Supply,

"Qd=Qs"


"1200-P = 120+3P"

Solving the above

"4P = 1080"

Hence Equilibrium price

"P = Rs 270"


When price Rises to Rs 400

"Qs = 120 + (3 \times400)"

"Qs = 1320"

and

"Qd = 1200 - 400 = 800"

From the above price Rd 400, we can say that There is more supply than Demand because the price is high.


When Price rises to Rs 120

"Qs = 120 + (3 \times120)"

"Qs = Rs 480"

And

"Qd = 1200 - 120 = Rs 1080"


The above price change implies that there is more demand than supply as the price is low.